OpenStudy (anonymous):

The stage is a right triangle with sides of a = 10 and b = 20 feet. How long is the other side if it is side c in this formula? (Use a^{2}+b ^{2}=c ^{2} and round your solution to the nearest tenth.) A. 25.6 ft B. 30.0 ft C. 22.4 ft D. 10 ft

OpenStudy (anonymous):

\[a^{2}+b ^{2}=c ^{2}\]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\large a^{2}+b^{2}=c^{2}\] \[\large 10^{2}+20^{2}=c^{2}\] \[\large 100+400=c^{2}\] ... ... ... \[\large c = ??\]

OpenStudy (anonymous):

\[500^{2}\]

jimthompson5910 (jim_thompson5910):

\[\large 100+400=c^{2}\] \[\large 500=c^{2}\] \[\large c^{2} = 500\] ... ... ... \[\large c = ??\]

OpenStudy (anonymous):

i dont understand what do we do ?

jimthompson5910 (jim_thompson5910):

example if c^2 = 36, then c = 6 since 6^2 = 36

jimthompson5910 (jim_thompson5910):

so if c^2 = 500, then c = ???

OpenStudy (anonymous):

25.6?

jimthompson5910 (jim_thompson5910):

take the square root of both sides

OpenStudy (anonymous):

thats the closest isnt it ?

OpenStudy (anonymous):

well 22.4

jimthompson5910 (jim_thompson5910):

22.4 is much closer